0850: Day starts: Drink 1L of water to hydrate body and counteract dehydration caused by long term exposure to dry, cool air conditioned laboratory.
0855: System checks. Inspected the ion gauge pressures, the status of all pumps and LN2 level. Inserted MIG into beam path in preparation for BEP flux measurements.
0856: Outgas Ga. Executed a program in the Ga cell’s PID controller than ramps up to 975°C in 30 minutes, holds for 30 minutes and then ramps down to 950°C and holds.
0900: As flux check. As shutter and valve were opened. Fluxes were taken at several valve positions consistent with previous days. Fluxes are recorded in the “flux check record” tab of the log spread sheet.
0915: As flux check completed. Gathered flux values were compared with legacy data for continuity checks. As shutter is closed and background As4 is pumped from the system. Vacuum recovery progresses prior to Ga flux check.
0920: Ga flux check. Ga shutter was opened for 60 seconds and the flux was recorded, averaged over 10 seconds. The shutter was then closed, the background flux was sampled, subtracted and the “flux check record” tab of the log spread sheet was updated. Current flux value is compared to legacy data for continuity.
0930: Colleague absent. In the absence of a colleague to discuss the strategy, the next sample was planned based my analysis of previous samples.
1000: Sample 1: Start. Transferred sample from outgas stage to growth chamber and manually began recipe.
1010: Sample 2 Outgas.
1155: Sample 1: Complete.
1200: Lunch and consumption of another litre of water.
1230: Sample 2: Start. Transferred sample from outgas stage to growth chamber and manually began recipe.
1420: Sample 2: Complete
1430: Sample 3: Considered. Theoretically a third sample could now be grown, however that would consume the same intended to grow first tomorrow. Incurring a 2 hour penalty and making it difficult to grow 2 samples tomorrow at the established 0.25 sample/hour rate. In order to avoid this problem another sample holder (platen) is required, making a total of 4. Two spare platens have been abrasively cleaned and are waiting etching, rinsing, baking and outgassing. The current platen 1 needs removing and cleaning as RHEED access is currently restricted.
1500: Sample load/unload. The two samples in the FEL (grown today) were removed and placed in labelled boxes. Two new Si(111) substrates were mounted and the FEL pumped down manually and the degas cycle started . The intention to implement “automatic FEL pump down and degas” was reaffirmed whilst shaking a fist at the heavens. Samples were passed to a colleague for SEM analysis.
1525: Update log. Sample locations were updated on the “locations” tab of the log spread sheet to reflect new sample locations and status.
1700: Day (almost) ends. Drink another litre of water. Go home.
1930: Analysis and presentation. Started preparing power point presentation slides for tomorrow’s weekly meeting. Analysis and comparison of results was very valuable. I noticed a new trend I had not seen before.
2130: Day (really) ends. Analysis and Presentation complete.
Total 10 hours. Growth: 2 samples. Samples/hour: 0.2.